What is the dot product of u = <-9, 4> and v = <1,3>? Thank you!

Answer by ikleyn(52776)

.
It is (-9)*1 + 4*3 = -9 + 12 = 3.

On dot-product see the lesson
    - Formula for Dot-product of vectors in a plane via the vectors components
in this site.

See also the lessons associated with it.


Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic
"Dot-product for vectors in a coordinate plane ".